OpenStudy (anonymous):

if y=3x-2, z=3y-2,w=3z-2, what is the value of w when x=2

OpenStudy (anonymous):

w=3z-2 z=3y-2 __________ w=3*(3y-2)-2 y=3x-3 ____________ w=3*(3*(3x-3)-2)-2 ________________ x=2 w=3*(3*(3*2-3)-2)-2 w=3*(3*(3-2)-2 w=3*(3*1-2) w=3*1 w=3